# DreamO: A Unified Framework for Image Customization

### v4

@@ 1. Introduction
Due to the high-quality image generation and stable performance of diffusion models (Ho et al., 2020), substantial research efforts focus on controllable generation by leveraging their generative priors (Mou et al., 2024; Zhang et al., 2023). Among these, image customization aims to ensure that generated outputs remain consistent with a reference image in specific attributes, such as identity (Wang et al., 2024a; Xiao et al., 2024b; Guo et al., 2024), object appearance (Huang et al., 2024a; Ruiz et al., 2023; He et al., 2025), virtual try-on (Choi et al., 2024; Luan et al., 2025; Wan et al., 2025), and style (Xing et al., 2024; Qi et al., 2024; Wu et al., 2021). Despite the abundance of task-specific approaches, developing a unified framework for image customization remains a challenge.
@@
Early research Composer (Huang et al., 2023) jointly trains a diffusion model with multi-condition input, e.g., depth, color, sketch. Some methods (Qin et al., 2023; Zhao et al., 2023) train additional control blocks (Zhang et al., 2023; Mou et al., 2024) to support general spatial control on the generation result, which greatly saves training costs. However, their control ability is restricted to some simple spatial conditions, and the interactions between conditions are rigid and have control redundancy. Recently, the DiT (Peebles and Xie, 2023) framework greatly scales up the performance of diffusion models. Based on DiT, OminiControl (Tan et al., 2024) proposes to integrate image conditions by unified sequence with diffusion latent. It can perform various customization tasks, e.g., identity, color, and layout. Despite its advantages, OminiControl is trained separately on different tasks, struggling to process multiple conditions. Recently, OmniGen (Xiao et al., 2024a) trains a general generation control based on a pre-trained large language model (Abdin et al., 2024) (LLMs). UniReal (Chen et al., 2024b) achieves this through video generation pretraining followed by full-model post-training. However, we argue that high-quality, multi-concept image customization cannot be achieved by merely leveraging the general capabilities of large language models like OmniGen or video models such as UniReal. Instead, it requires specialized architectural designs. Currently, the research community lacks an efficient and effective method specifically tailored for image customization under multi-concept and multi-conditional scenarios.
@@
In this paper, we design a unified image customization approach based on a pre-trained DiT. With a low training cost and a single model, our method can support various conditions (e.g., identity, subject, try-on, and style) and enables interactions among different kinds of condition, as shown in 1. Specifically, we follow the unified sequence conditioning format introduced in OminiControl (Tan et al., 2024), and introduce a routing constraint on the internal representations of DiT during training. This ensures content fidelity and promotes the disentanglement of different control conditions. A placeholder strategy is also designed to enable control over the placement of conditions in the generated results. In addition, we construct large-scale training data covering multiple tasks and design a progressive training strategy. This enables the model to progressively acquire robust and generalized image customization capabilities.
@@ 2.1. Diffusion Models
As a powerful generation paradigm, the diffusion model (Ho et al., 2020; Dhariwal and Nichol, 2021) rapidly dominates the image generation community. Its high generation quality and stable performance have been successfully applied to various tasks , e.g., text-to-image generation (Saharia et al., 2022a; Ramesh et al., 2022; Nichol et al., 2022; Rombach et al., 2022) and image editing (Hertz et al., 2022; Meng et al., 2021; Yang et al., 2023; Mou et al., 2023). Some strategies, such as latent diffusion (Rombach et al., 2022) and flow matching sampling (Lipman et al., 2022), are proposed to enhance the performance. Most early works (Rombach et al., 2022; Nichol et al., 2022) utilize the UNet architecture as the diffusion model. Recently, the diffusion transformer (Peebles and Xie, 2023) (DiT) architecture has emerged as a superior choice, offering improved performance through straightforward scalability.
@@ 2.2. Controllable Image Generation
Advancements in diffusion models drive the rapid development of controllable image generation. In this community, text is the most fundamental conditioning input, e.g., Stable Diffusion (Rombach et al., 2022), Imagen (Saharia et al., 2022b), and DALL-E2 (Ramesh et al., 2022). To achieve accurate spatial control, some methods, e.g., ControlNet (Zhang et al., 2023) and T2I-Adapter (Mou et al., 2024), propose adding control modules on pre-trained diffusion models. UniControl (Qin et al., 2023; Zhao et al., 2023) propose to unify different spatial conditions with a joint condition input. In addition to spatial control, some unspatial conditions are also studied. The IP-Adapter (Ye et al., 2023) utilizes cross-attention to inject image prompt in the diffusion model to control some unspatial properties, e.g., identity and style. In addition to these representative works, other related attempts (Gal et al., 2022; He et al., 2025; Hua et al., 2023; Kumari et al., 2023; Li et al., 2023; Ma et al., 2024a, b) also help to broaden the scope of controllable image generation.
Recent advancements in DiT-based diffusion models further promote the development of controllable image generation. For instance, in-context LoRA (Huang et al., 2024b) and OminiControl (Tan et al., 2024) introduce a novel approach by concatenating all input tokens (i.e., text, image, and conditions) and training LoRA with task-specific datasets for various applications. Subsequently, OmniGen (Xiao et al., 2024a) and UniReal (Chen et al., 2024b) optimize the entire diffusion model in multiple stages on larger-scale training data, achieving improved understanding of input conditions.
@@ 2.3. Cross-attention Routing in Diffusion Models
Existing studies (e.g., Prompt-to-Prompt (Hertz et al., 2022)) demonstrate that text-visual cross-attention maps inherently establish spatial-semantic correspondence between linguistic tokens and visual generation, i.e., the similarity response aligns with the spatial region of the corresponding subject in the generated result. Building upon this observation, UniPortrait (He et al., 2024) constrains the influence region of condition features for identity-specific generation in multi-face scenarios, and AnyStory (He et al., 2025) further extends this approach to subject-driven generation. Some recent works (Yamaguchi and Yanai, 2024) show that the cross-attention map in the DiT framework also exhibits spatial properties. In this paper, we explore routing constraints in the DiT framework.
@@ 3.1. Preliminaries
The Diffusion Transformer (DiT) model (Peebles and Xie, 2023) employs a transformer as the denoising network to refine diffusion latent. Specifically, in the input, the 2D image latent $\mathbf{z}_{t}\in\mathbb{R}^{c\times w\times h}$ is patchified into a sequence of 1D tokens $\mathbf{z}_{t}\in\mathbb{R}^{c\times(\frac{w}{p}\times\frac{h}{p})}$, where $(w,h)$ is spatial size, $c$ is the number of channels, and $p$ is the patch size. The image and text tokens are concatenated and processed by the DiT model in a unified manner. Apart from model architectures, more efficient sampling strategies (e.g., Flow Matching (Lipman et al., 2022)) are also proposed. Unlike DDPM (Ho et al., 2020), Flow Matching conducts the forward process by linearly interpolating between noise and data in a straight line. At the time step $t$, latent $\mathbf{z}_{t}$ is defined as: $\mathbf{z}_{t}=(1-t)\mathbf{z}_{0}+t\epsilon$, where $\mathbf{z}_{0}$ is the clean image, and $\epsilon\in\mathcal{N}(0,1)$ is the Gaussian noise. The model is trained to directly regress the target velocity given the noised latent $\mathbf{z}_{t}$, timestep t, and condition $y$. The objective is to minimize the mean squared error:
@@
where $\mathcal{V}_{\theta}$ refers to the diffusion model. The DiT framework and Flow Matching are widely used in some recent diffusion models, such as Stable Diffusion3 (Esser et al., 2024) and Flux (flu, 2023a).
@@ 3.2. Overview
An overview of our method is presented in Fig. 2. Specifically, we utilize the Flux-1.0-dev (flu, 2023a) as the base model to build a unified framework for image customization, e.g., style, identity, subject appearance, and try-on. Given $n$ condition images $\mathbf{C}=\{\mathbf{C}_{1},...,\mathbf{C}_{n}\}$, we first reuse the VAE (Kingma et al., 2013) of Flux to encode the condition image to the same latent space as noisy latent. Note that the size of the condition image is flexible. Higher resolutions are recommended for detail-rich images to preserve clarity, while lower resolutions are sufficient for images with fewer details, thereby reducing compression costs. Then, all tokens (i.e., image, text, condition) are concatenated along the sequence dimension and fed into Flux. To enable the model to incorporate the condition input, we introduce a condition mapping layer at the input of Flux. The position embedding (PE) of condition tokens is aligned with that of the noisy latent using Rotary Position Embedding (RoPE) (Su et al., 2024). Inspired by non-overlapping position embedding in OminiControl (Tan et al., 2024), we extend these embeddings along the diagonal in a similar fashion. In addition, we introduce a trainable and index-wise condition embedding (CE) $\mathbb{R}^{10\times c}$, which is directly added to condition tokens. Following OminiControl (Tan et al., 2024), we integrate Low-Rank Adaptation (LoRA) (Ryu, 2023) modules into Flux as trainable parameters.
@@ 3.3. Routing Constraint
Inspired by UniPortrait (He et al., 2024) and AnyStory (He et al., 2025), in this paper, we design routing constraint in the DiT framework for general image customization tasks. As illustrated in Fig. 2, within the condition-guided framework, cross-attention exists between condition images and the generation result:
@@ 3.4. Training Data Construction
Identity paired data. Since high-quality identity paired data (Li et al., 2024) is difficult to collect from the Internet, we adopt the open-source ID customization method PuLID (Guo et al., 2024) for dataset construction. Specifically, we generate two images of the same identity using PuLID-FLUX, which then serve as mutual references. We also provide PuLID-SDXL with a reference face image and a text prompt describing the desired style to produce stylized training pairs. Finally, we collect 150K photorealistic data and 60K stylized identity data.
Subject-driven data. For single-subject-driven image customization, we utilize the Subject200K (Tan et al., 2024) dataset as part of training data. To rectify the absence of character-related conditions, we collect 100K paired character-related data through retrieval. For multi-subject-driven image customization, we construct some two-column images through concatenation on the Subject200K dataset. In addition, we employ X2I-subject (Xiao et al., 2024a) dataset in multi-subject-driven training. To improve human-driven generation, we develop a pipeline similar to MovieGen (Polyak et al., 2025). Starting with a long-video dataset, we apply content-aware scene detection to extract short clips. Mask2Former (Cheng et al., 2022) is used to generate human masks for key frames and perform object tracking. For cross-clip instance matching, we use SigLip (Zhai et al., 2023) embeddings and clustering.
Try-on data. We create a paired try-on dataset from two sources. One part consists of paired model and clothing images collected directly from the Web. For the other, we first crawl high-quality model images as ground truth, then use image segmentation (Jin, 2023; Jin et al., 2024) to extract clothing items and generate the corresponding pairs. All images are manually filtered to remove low-quality samples, resulting in a dataset of 500K try-on pairs.
Style-driven data. This paper tackles two style transfer tasks: (1) style reference control with text descriptions of content, and (2) style reference with content reference images. For the first task, we use an internal style customization model based on SDXL to generate images with the same style but different content from two distinct prompts. For the second task, training requires style reference images, content reference images, and target images, where the target image must match both the style and content structure of the reference images. Based on the type 1 dataset, we produce the content reference for each style image using Canny-guided Flux (flu, 2023c). More details are provided in the Appendix.
Routing mask extraction. To obtain the labels for the routing constraint (i.e., Eq.3), we use LISA(Lai et al., 2024) to extract mask of object conditioned on the text descriptions. In certain complex datasets, we employ InternVL (Chen et al., 2024a) to generate description of the target object. More details are provided in the Appendix.
@@ 3.5. Progressive Training Process
In experiments, we find that training directly on all data makes convergence difficult. This is mainly due to the limited capacity of LoRA (Ryu, 2023) optimization, making it difficult for the model to capture task-specific capabilities under complex data distributions. In addition to convergence, the quality of the output after training deviates from the original prior of Flux (flu, 2023a). This divergence is caused by the impact of some low-quality training samples.
@@
To address these issues, we design a progressive training strategy that allows the model to smoothly converge across different tasks while mitigating the influence of training data on the generation prior of Flux. The training pipeline is shown in Fig. 4. Specifically, we first optimize the model on subject-driven training data to initiate the model with a consistency-preserving capability. Note that the Subject200K (Tan et al., 2024) training data is generated by the base model (i.e., Flux), thus shares a similar distribution with the model generation space, which facilitates fast convergence. Since the X2I-subject (Xiao et al., 2024a) dataset is synthetically generated by MS-Diffusion (Wang et al., 2024b), a lot of training samples contain undesired artifacts and distortions. Therefore, during this warm-up stage, the two-column Subject200K images described in Sec. 3.4 are also utilized as part of the training data to facilitate rapid convergence of the multi-reference control. The right part of Fig. 4 illustrates that after the first training stage, the model acquires an initial subject-driven generation capability and presents strong text-following performance. In the second training stage, we incorporate all the training data and perform full-data tuning. This allows the model to further converge on all subtasks defined in this work.
@@ 4.1. Implementation Details
In this paper, we adopt Flux-1.0-dev as the base model. The rank of LoRA (Ryu, 2023) is set as 128, resulting in a total parameter increase of 478M. During training, we employ the Adam (Loshchilov and Hutter, 2017) optimizer with a learning rate of 4e-5 and train on 8 NVIDIA A100 80G GPUs. The batch size is set as 8. The first training stage consists of 20K iterations, followed by 90K iterations in the second stage, and finally 3K iterations in the last training stage. In inference, we use Flux-Turbo (flu, 2023b) for acceleration, enabling the generation of $1024\times 1024$ results within 10s. Unless specified, all results in this paper are based on the Turbo model.

### v5

@@ 1. Introduction
Due to the high-quality image generation and stable performance of diffusion models (ddpm), substantial research efforts focus on controllable generation by leveraging their generative priors (t2iadapter; controlnet). Among these, image customization aims to ensure that generated outputs remain consistent with a reference image in specific attributes, such as identity (instantid; fastcomposer; pulid), object appearance (realcustom; dreambooth; anystory), virtual try-on (tryon-1; tryon-2; tryon-3), and style (csgo; deadiff; stylealign). Despite the abundance of task-specific approaches, developing a unified framework for image customization remains a challenge.
@@
Early research Composer (composer) jointly trains a diffusion model with multi-condition input, e.g., depth, color, sketch. Some methods (unicontrol-1; unicontrol-2) train additional control blocks (controlnet; t2iadapter) to support general spatial control on the generation result, which greatly saves training costs. However, their control ability is restricted to some simple spatial conditions, and the interactions between conditions are rigid and have control redundancy. Recently, the DiT (dit) framework greatly scales up the performance of diffusion models. Based on DiT, OminiControl (ominicontrol) proposes to integrate image conditions by unified sequence with diffusion latent. It can perform various customization tasks, e.g., identity, color, and layout. Despite its advantages, OminiControl is trained separately on different tasks, struggling to process multiple conditions. Recently, OmniGen (omnigen) trains a general generation control based on a pre-trained large language model (phi3) (LLMs). UniReal (unireal) achieves this through video generation pretraining followed by full-model post-training. However, we argue that high-quality, multi-concept image customization cannot be achieved by merely leveraging the general capabilities of large language models like OmniGen or video models such as UniReal. Instead, it requires specialized architectural designs. Currently, the research community lacks an efficient and effective method specifically tailored for image customization under multi-concept and multi-conditional scenarios.
@@
In this paper, we design a unified image customization approach based on a pre-trained DiT. With a low training cost and a single model, our method can support various conditions (e.g., identity, subject, try-on, and style) and enables interactions among different kinds of condition, as shown in 1. Specifically, we follow the unified sequence conditioning format introduced in OminiControl (ominicontrol), and introduce a routing constraint on the internal representations of DiT during training. This ensures content fidelity and promotes the disentanglement of different control conditions. A placeholder strategy is also designed to enable control over the placement of conditions in the generated results. In addition, we construct large-scale training data covering multiple tasks and design a progressive training strategy. This enables the model to progressively acquire robust and generalized image customization capabilities.
@@ 2.1. Diffusion Models
As a powerful generation paradigm, the diffusion model (ddpm; diff_beat) rapidly dominates the image generation community. Its high generation quality and stable performance have been successfully applied to various tasks , e.g., text-to-image generation (t2i1; dall-e2; glid; ldm) and image editing (p2p; sdedit; paint-by; dragondiffusion). Some strategies, such as latent diffusion (ldm) and flow matching sampling (flow), are proposed to enhance the performance. Most early works (ldm; glid) utilize the UNet architecture as the diffusion model. Recently, the diffusion transformer (dit) (DiT) architecture has emerged as a superior choice, offering improved performance through straightforward scalability.
@@ 2.2. Controllable Image Generation
Advancements in diffusion models drive the rapid development of controllable image generation. In this community, text is the most fundamental conditioning input, e.g., Stable Diffusion (ldm), Imagen (imagen), and DALL-E2 (dall-e2). To achieve accurate spatial control, some methods, e.g., ControlNet (controlnet) and T2I-Adapter (t2iadapter), propose adding control modules on pre-trained diffusion models. UniControl (unicontrol-1; unicontrol-2) propose to unify different spatial conditions with a joint condition input. In addition to spatial control, some unspatial conditions are also studied. The IP-Adapter (ipadapter) utilizes cross-attention to inject image prompt in the diffusion model to control some unspatial properties, e.g., identity and style. In addition to these representative works, other related attempts (relat_1; anystory; relat_3; relat_4; relat_5; relat_6; relat_7) also help to broaden the scope of controllable image generation.
Recent advancements in DiT-based diffusion models further promote the development of controllable image generation. For instance, in-context LoRA (Incontext) and OminiControl (ominicontrol) introduce a novel approach by concatenating all input tokens (i.e., text, image, and conditions) and training LoRA with task-specific datasets for various applications. Subsequently, OmniGen (omnigen) and UniReal (unireal) optimize the entire diffusion model in multiple stages on larger-scale training data, achieving improved understanding of input conditions.
@@ 2.3. Cross-attention Routing in Diffusion Models
Existing studies (e.g., Prompt-to-Prompt (p2p)) demonstrate that text-visual cross-attention maps inherently establish spatial-semantic correspondence between linguistic tokens and visual generation, i.e., the similarity response aligns with the spatial region of the corresponding subject in the generated result. Building upon this observation, UniPortrait (uniportrait) constrains the influence region of condition features for identity-specific generation in multi-face scenarios, and AnyStory (anystory) further extends this approach to subject-driven generation. Some recent works (seg_att) show that the cross-attention map in the DiT framework also exhibits spatial properties. In this paper, we explore routing constraints in the DiT framework.
@@ 3.1. Preliminaries
The Diffusion Transformer (DiT) model (dit) employs a transformer as the denoising network to refine diffusion latent. Specifically, in the input, the 2D image latent $\mathbf{z}_{t}\in\mathbb{R}^{c\times w\times h}$ is patchified into a sequence of 1D tokens $\mathbf{z}_{t}\in\mathbb{R}^{c\times(\frac{w}{p}\times\frac{h}{p})}$, where $(w,h)$ is spatial size, $c$ is the number of channels, and $p$ is the patch size. The image and text tokens are concatenated and processed by the DiT model in a unified manner. Apart from model architectures, more efficient sampling strategies (e.g., Flow Matching (flow)) are also proposed. Unlike DDPM (ddpm), Flow Matching conducts the forward process by linearly interpolating between noise and data in a straight line. At the time step $t$, latent $\mathbf{z}_{t}$ is defined as: $\mathbf{z}_{t}=(1-t)\mathbf{z}_{0}+t\epsilon$, where $\mathbf{z}_{0}$ is the clean image, and $\epsilon\in\mathcal{N}(0,1)$ is the Gaussian noise. The model is trained to directly regress the target velocity given the noised latent $\mathbf{z}_{t}$, timestep t, and condition $y$. The objective is to minimize the mean squared error:
@@
where $\mathcal{V}_{\theta}$ refers to the diffusion model. The DiT framework and Flow Matching are widely used in some recent diffusion models, such as Stable Diffusion3 (sd3) and Flux (flux).
@@ 3.2. Overview
An overview of our method is presented in Fig. 2. Specifically, we utilize the Flux-1.0-dev (flux) as the base model to build a unified framework for image customization, e.g., style, identity, subject appearance, and try-on. Given $n$ condition images $\mathbf{C}=\{\mathbf{C}_{1},...,\mathbf{C}_{n}\}$, we first reuse the VAE (vae) of Flux to encode the condition image to the same latent space as noisy latent. Note that the size of the condition image is flexible. Higher resolutions are recommended for detail-rich images to preserve clarity, while lower resolutions are sufficient for images with fewer details, thereby reducing compression costs. Then, all tokens (i.e., image, text, condition) are concatenated along the sequence dimension and fed into Flux. To enable the model to incorporate the condition input, we introduce a condition mapping layer at the input of Flux. The position embedding (PE) of condition tokens is aligned with that of the noisy latent using Rotary Position Embedding (RoPE) (rope). Inspired by non-overlapping position embedding in OminiControl (ominicontrol), we extend these embeddings along the diagonal in a similar fashion. In addition, we introduce a trainable and index-wise condition embedding (CE) $\mathbb{R}^{10\times c}$, which is directly added to condition tokens. Following OminiControl (ominicontrol), we integrate Low-Rank Adaptation (LoRA) (lora) modules into Flux as trainable parameters.
@@ 3.3. Routing Constraint
Inspired by UniPortrait (uniportrait) and AnyStory (anystory), in this paper, we design routing constraint in the DiT framework for general image customization tasks. As illustrated in Fig. 2, within the condition-guided framework, cross-attention exists between condition images and the generation result:
@@ 3.4. Training Data Construction
Identity paired data. Since high-quality identity paired data (photomaker) is difficult to collect from the Internet, we adopt the open-source ID customization method PuLID (pulid) for dataset construction. Specifically, we generate two images of the same identity using PuLID-FLUX, which then serve as mutual references. We also provide PuLID-SDXL with a reference face image and a text prompt describing the desired style to produce stylized training pairs. Finally, we collect 150K photorealistic data and 60K stylized identity data.
Subject-driven data. For single-subject-driven image customization, we utilize the Subject200K (ominicontrol) dataset as part of training data. To rectify the absence of character-related conditions, we collect 100K paired character-related data through retrieval. For multi-subject-driven image customization, we construct some two-column images through concatenation on the Subject200K dataset. In addition, we employ X2I-subject (omnigen) dataset in multi-subject-driven training. To improve human-driven generation, we develop a pipeline similar to MovieGen (moive_gen). Starting with a long-video dataset, we apply content-aware scene detection to extract short clips. Mask2Former (mask2former) is used to generate human masks for key frames and perform object tracking. For cross-clip instance matching, we use SigLip (zhai2023sigmoid) embeddings and clustering.
Try-on data. We create a paired try-on dataset from two sources. One part consists of paired model and clothing images collected directly from the Web. For the other, we first crawl high-quality model images as ground truth, then use image segmentation (jin2023sssegmenation; jin2024idrnet) to extract clothing items and generate the corresponding pairs. All images are manually filtered to remove low-quality samples, resulting in a dataset of 500K try-on pairs.
Style-driven data. This paper tackles two style transfer tasks: (1) style reference control with text descriptions of content, and (2) style reference with content reference images. For the first task, we use an internal style customization model based on SDXL to generate images with the same style but different content from two distinct prompts. For the second task, training requires style reference images, content reference images, and target images, where the target image must match both the style and content structure of the reference images. Based on the type 1 dataset, we produce the content reference for each style image using Canny-guided Flux (flux_canny). More details are provided in the Appendix.
Routing mask extraction. To obtain the labels for the routing constraint (i.e., Eq.3), we use LISA(lisa) to extract mask of object conditioned on the text descriptions. In certain complex datasets, we employ InternVL (internvl) to generate description of the target object. More details are provided in the Appendix.
@@ 3.5. Progressive Training Process
In experiments, we find that training directly on all data makes convergence difficult. This is mainly due to the limited capacity of LoRA (lora) optimization, making it difficult for the model to capture task-specific capabilities under complex data distributions. In addition to convergence, the quality of the output after training deviates from the original prior of Flux (flux). This divergence is caused by the impact of some low-quality training samples.
@@
To address these issues, we design a progressive training strategy that allows the model to smoothly converge across different tasks while mitigating the influence of training data on the generation prior of Flux. The training pipeline is shown in Fig. 4. Specifically, we first optimize the model on subject-driven training data to initiate the model with a consistency-preserving capability. Note that the Subject200K (ominicontrol) training data is generated by the base model (i.e., Flux), thus shares a similar distribution with the model generation space, which facilitates fast convergence. Since the X2I-subject (omnigen) dataset is synthetically generated by MS-Diffusion (ms_diffusion), a lot of training samples contain undesired artifacts and distortions. Therefore, during this warm-up stage, the two-column Subject200K images described in Sec. 3.4 are also utilized as part of the training data to facilitate rapid convergence of the multi-reference control. The right part of Fig. 4 illustrates that after the first training stage, the model acquires an initial subject-driven generation capability and presents strong text-following performance. In the second training stage, we incorporate all the training data and perform full-data tuning. This allows the model to further converge on all subtasks defined in this work.
@@ 4.1. Implementation Details
In this paper, we adopt Flux-1.0-dev as the base model. The rank of LoRA (lora) is set as 128, resulting in a total parameter increase of 478M. During training, we employ the Adam (adam) optimizer with a learning rate of 4e-5 and train on 8 NVIDIA A100 80G GPUs. The batch size is set as 8. The first training stage consists of 20K iterations, followed by 90K iterations in the second stage, and finally 3K iterations in the last training stage. In inference, we use Flux-Turbo (flux_turbo) for acceleration, enabling the generation of $1024\times 1024$ results within 10s. Unless specified, all results in this paper are based on the Turbo model. Some of the example inputs are processed using BEN2 (ben2) to remove background.